# FineReason: Evaluating and Improving LLMs’ Deliberate Reasoning through Reflective Puzzle Solving

### v1

@@ Abstract
Many challenging reasoning tasks require not just rapid, intuitive responses, but a more deliberate, multi-step approach. Recent progress in large language models (LLMs) highlights an important shift from the “System 1” way of quick reactions to the “System 2” style of reflection-and-correction problem solving. However, current benchmarks heavily rely on the final-answer accuracy, leaving much of a model’s intermediate reasoning steps unexamined. This fails to assess the model’s ability to reflect and rectify mistakes within the reasoning process. To bridge this gap, we introduce FineReason, a logic-puzzle benchmark for fine-grained evaluation of LLMs’ reasoning capabilities. Each puzzle can be decomposed into atomic steps, making it ideal for rigorous validation of intermediate correctness. Building on this, we introduce two tasks: state checking, and state transition, for a comprehensive evaluation of how models assess the current situation and plan the next move. To support broader research, we also provide a puzzle training set aimed at enhancing performance on general mathematical tasks. We show that models trained on our state checking and transition data demonstrate gains in math reasoning by up to $5.1\%$ on GSM8K. Our data and code can be found at https://github.com/DAMO-NLP-SG/FineReason.
@@
Guizhen Chen1,2thanks: Guizhen is under the Joint PhD Program between Alibaba and NTU.  Weiwen Xu2  Hao Zhang2  Hou Pong Chan2  Chaoqun Liu2 Lidong Bing2Deli Zhao2Anh Tuan Luu1Yu Rong2 1 Nanyang Technological University, Singapore 2 DAMO Academy, Alibaba Group, Singapore
@@ 1 Introduction
Figure 1: An illustration of stepwise state checking and transition in a Sudoku solution tree.
Figure 2: An illustration of four puzzle categories in FineReason.
@@
In cognitive science, human reasoning is typically characterized by two distinct systems: (i) System 1, which is fast, automatic, and effortless, and (ii) System 2, which is slow, analytical, and effortful (Kahneman, 2011). System 2 reasoning enables humans to proactively anticipate future outcomes, reassess intermediate states, and iteratively refine solutions Yao et al. (2023), thereby allowing them to tackle more complex reasoning tasks. Recent studies suggest that large language models (LLMs) can attain advantages akin to those of System 2 reasoning OpenAI (2024); Snell et al. (2024); Guo et al. (2025); Team (2025). Instead of generating an answer directly, these models can engage in iterative reflection and correction to refine their reasoning processes Shinn et al. (2023), leading to impressive performance on reasoning-intensive tasks such as mathematics and coding Qin et al. (2024); Muennighoff et al. (2025).
@@
Despite these improvements, the underlying reasoning mechanisms responsible for these enhancements remain underexplored. Existing reasoning benchmarks primarily focus on the final-answer accuracy Hendrycks et al. (2020); Cobbe et al. (2021); Chen et al. (2021); Hendrycks et al. (2021), which offers limited insight into LLMs’ internal reasoning processes, such as reflection and correction. For instance, a model might reach a correct conclusion through flawed reasoning Zelikman et al. (2022); Creswell et al. (2023); Lightman et al. (2024). This diminishes the trustworthiness of model outputs, posing potential risks in practical usage. Moreover, models may achieve high accuracy by exploiting superficial patterns in the training data Roelofs et al. (2019); Enström et al. (2024), making it difficult to ascertain whether the observed performance gain truly stems from deliberate reasoning. Therefore, there is a pressing need for more comprehensive reasoning benchmarks that assess the integrity of intermediate processes.
@@
In this work, we present FineReason, a logic-puzzle benchmark for granular evaluation of LLMs’ reasoning capabilities. FineReason includes four types of puzzles: Sudoku, Graph Coloring, Game of 24, and Grid Puzzles. Solving a logic puzzle involves a series of discrete steps, and its explicit rules make it straightforward for validating the intermediate states. To structure our evaluation, we propose two key actions in each atomic step, state checking and state transition, as shown in Figure 1. State checking predicts whether the current state can lead to a solvable solution Agarwal et al. (2019); Wang et al. (2024). It captures both retrospective evaluation of prior steps Lightman et al. (2024) and prospective analysis of future steps. Meanwhile, state transition focuses on determining the next valid step, either moving forward or backtracking to the previous state. Together, these two tasks cover the entire puzzle-solving process, revealing the internal reasoning processes of reflection, correction, and exploration of alternative paths in large language models.
Our evaluation results show that OpenAI-o1 OpenAI (2024) outperforms Gemini-2.0-Flash-Thinking Google (2024) by a large margin of 19.7%, despite both being reasoning-oriented models that already excel at common math and coding tasks. On the other hand, general-purpose models, such as GPT-4o OpenAI et al. (2024), GPT-3.5 OpenAI (2022), Qwen2.5-72b-Instruct Qwen et al. (2025), and Gemini-2.0-Flash Google (2024) struggle with deep reasoning, often making near-random guesses during state checking and showing extremely low performance in state transition task. To improve the reasoning capabilities of models, we develop a specialized training set on state checking and transition. Integrating our dataset with math training data consistently boosts performance in mathematical reasoning. For example, when applied to GSM8K using the DeepSeek-R1-Distill-Qwen-7B model, our state checking and transition data improve math reasoning accuracy from 82.3% to 87.4%, compared to models trained exclusively on math data. This suggests that our data offers a straightforward way to boost the reasoning abilities of LLMs, thereby enhancing performance on other reasoning-intensive tasks.
@@
Our main contributions are three-fold: 1) We introduce FineReason, a fine-grained, puzzle-based benchmark accompanied by systematic evaluations on state checking and transition, to comprehensively evaluate reasoning capabilities of LLMs. 2) Our experiments reveal substantial limitations in deep reasoning among general-purpose models, and even in the leading reasoning model, Gemini-2.0-Flash-Thinking, leaving substantial room for improvement. 3) We develop a supplementary training set focused on puzzles, which enhances general mathematical reasoning in LLMs.
@@ 2 FineReason
We present FineReason, a logic-puzzle benchmark that comprehensively assesses LLMs’ reasoning abilities through stepwise evaluation of state checking and transition. An overview of each puzzle is shown in Figure 2. Inspired by the adage “think twice before acting,” the two actions capture how models assess the current situation (i.e., state checking) and plan the next move (i.e., state transition), which is essential for deliberate reasoning.
Formally, we represent the reasoning process of a logic puzzle as $p=\{p_{1},p_{2},...,p_{n}\}$, where $n$ denotes the number of atomic steps. Each step $p_{i}$ consists of a puzzle state $s_{i}$ and two actions: state checking $a^{c}_{i}$ and state transition $a^{t}_{i}$, i.e., $p_{i}=(s_{i},a^{c}_{i},a^{t}_{i})$. Applying these actions to $s_{i}$ produces the next state $s_{i+1}$. The puzzle-solving process begins at an initial state $s_{1}$ and proceeds through a sequence of such atomic steps until reaching the solution state $s_{n}$ that satisfies all constraints.
In the following section, we first introduce a tree-based approach for step decomposition(§2.1) in our puzzles. Next, we describe the two key actions – state checking and state transition – that facilitate fine-grained reasoning evaluation of models (§2.2).
@@ 2.1 Tree-based Puzzle Decomposition
Puzzle solving can be represented as a tree, where nodes correspond to intermediate states describing the completion stage of the current puzzle, and edges represent the execution of state checking and state transition, as illustrated in Figure 1. Edges are bidirectional, enabling both the exploration of child states and backtracking to the parent state when encountering dead ends. This process begins at the root node $s_{1}$ and terminates at a solution leaf $s_{n}$, potentially requiring multiple backtracks to explore alternate paths.
@@
Given a partially filled $9\times 9$ Sudoku grid, the aim is to fill the remaining cells such that each row, each column, and each of the nine $3\times 3$ subgrids contains all digits from 1 to 9 exactly once. A Sudoku state can be either a partially or fully completed $9\times 9$ grid. The former refers to any intermediate state $s_{i}$ encountered during the reasoning process, while the latter is the final solvable state $s_{n}$. A minimal move consists of either adding a number to the grid for exploration or removing a number for backtracking. Our Sudoku questions are collected from Kaggle. [^1] https://www.kaggle.com/datasets/bryanpark/sudoku
@@
The puzzle of graph coloring is to assign colors to each vertex in a graph such that no two adjacent vertices share the same color. Each puzzle specifies a maximum number of colors allowed in a graph. A graph coloring state is either a partially colored graph, denoted as $s_{i}$, or a completely colored graph, denoted as $s_{n}$. A minimal move involves either assigning a color to a vertex or removing a color from a vertex. To create the questions, we generate random graphs and find their respective chromatic numbers using the backtracking algorithm van Beek (2006).
@@
In Game of 24, given four numbers, the task is to apply basic arithmetic operations (addition, subtraction, multiplication, and division) in any order to reach exactly the value of 24. Each number must be used exactly once. Each state is a partial or complete arithmetic expression. The minimal move is to apply or unapply an operation to two of the remaining numbers. We use the dataset from Yao et al. (2023).
@@ 2.2 Evaluation Tasks
We define two key actions, state checking and state transition, which enable movement between states.
@@
State checking involves assessing if a given state $s_{i}$ can lead to a solvable leaf $s_{n}$. Based on our constructed puzzle tree, we label a state as solvable if at least one valid solution exists in the subtree of $s_{i}$. To ensure a diverse difficulty range, we uniformly sample both solvable and unsolvable states across different tree depths. For each sampled state, we prompt models to evaluate its solvability with puzzle rules, prior visited states, and the current state (see Appendix A.2). In general, state checking evaluates two key aspects: 1) It tests if a model can reflect on prior steps (i.e., states and actions) to avoid rule violations (e.g., preventing duplicate "1"s in a Sudoku row). 2) It requires models to anticipate potential dead ends by looking ahead. The second aspect, however, requires a higher level of reasoning to proactively detect states that are unsolvable.
@@
State transition involves making the minimal move defined in §2.1, which requires models to determine the next valid state. Based on the state-checking results, models should proceed if the state is solvable and backtrack otherwise. Specifically, at a solvable state, a correct transition would be an unvisited child of the given state. At an unsolvable state, the correct move is to backtrack to its parent state. To isolate the impact of state transition from incorrect state checking, our evaluation provides the ground-truth state-checking labels. We sample states from the puzzle tree and construct prompts with puzzle rules, prior visited states, the sampled state, and some unsolvable child states (see Appendix A.2). The inclusion of unsolvable child states is to assess whether models can effectively bypass these bad states.
@@
Table 2: A preliminary study of End-to-end puzzle-solving performance of LLMs.
@@ 3 Experimental Setup
Implementations.
To elicit the reasoning capability of LLMs, we leverage the zero-shot chain of thought (CoT) prompting Kojima et al. (2022). To ensure a genuine evaluation of LLM’s inherent reasoning capabilities, we explicitly include the instruction “Do not solve using programming” in the prompt, restricting the models from relying on programmatic solutions.
@@
We select the best-performing open and closed-source models, including 1) reasoning-oriented models with deep thinking: o1 OpenAI (2024), Gemini-2.0-Flash-Thinking (Gemini-FT, Google (2024)), and 2) general-purposed models that perform straightforward execution: GPT-4o OpenAI et al. (2024), GPT-3.5 OpenAI (2022), Gemini-2.0-Flash (Gemini-F, Google (2024)), and Qwen2.5-72B-Instruct (Qwen2.5-Inst, Qwen et al. (2025)).
@@ 4 Evaluation Results
In this section, we first present a preliminary evaluation of LLMs on end-to-end puzzle solving in §4.1, which reveals inconsistencies in model performance. To gain deeper insights beyond their end-to-end performance, we present our main results focusing on state checking and transition tasks in §4.2, followed by detailed analyses on models behaviors, error patterns and performance across different difficulty levels in the remaining section.
@@ 4.1 Preliminary: End-to-End Puzzle Solving
Table 2 presents an initial evaluation of LLMs on end-to-end puzzle-solving tasks. Despite their strong performance on coding and math tasks Qwen et al. (2025), these models are notably weak in end-to-end puzzle solving. Additionally, there are some counter-intuitive observations: Gemini-F outperforms Gemini-FT on Sudoku and Game of 24, yet struggle on the other two puzzles. These inconsistencies suggest that end-to-end puzzle solving alone may not be a reliable metric for assessing LLMs’ reasoning, emphasizing the need for more granular evaluation methods.
@@ 4.2 Main results
To explore models’ reasoning capabilities in greater depth, we present our evaluation results on state checking and transition in Table 3. The results reveal noticeable performance gaps between reasoning-oriented and general-purposed models across both tasks. On the state-checking task, reasoning-oriented models (o1 and Gemini-FT) consistently lead the performance in every puzzle category. In contrast, general-purposed models barely match or slightly surpass the random baseline in puzzles like Sudoku and Grid Puzzles. A similar trend is observed on the state-transition task, particularly for the Game of 24 and Grid Puzzles. These findings further support the view that inference-time scaling can substantially boost reasoning capabilities Snell et al. (2024); Muennighoff et al. (2025).
Among the reasoning models, Gemini-FT performs on par with o1 in state checking but consistently lags behind in state transition across all puzzle categories. This reveals weaknesses in Gemini-FT’s reasoning process, particularly in error revision. These findings align well with our practical experience using these models, which provides empirical evidence that FineReason offers a more accurate reflection of LLMs’ reasoning.
Moreover, we observe that most models exhibit a significant “execution gap”: they perform better in state checking than in state transition. This indicates a tendency for models to skip correct intermediate steps or find the short-cut to final answer rather than methodically progressing toward a solution.
4.3 State-Checking Patterns
A key requirement of state checking is to look ahead at potential paths and detect unsolvable states. To analyze models’ behaviors in this task, we report the state-checking precision, recall, and F1 scores in Table 4. We designate unsolvable states as positive cases. Therefore, recall measures how likely a model detects the dead ends, whereas precision shows the reliability of predicting unsolvable states.
We find that reasoning models can generally detect unsolvable states well, as indicated by the high F1 scores. As for general models (GPT-4o, Gemini-F, Qwen2.5-Inst), the recall is generally low in Sudoku but not in Game of 24. This difference could be attributed to a deeper puzzle tree of Sudoku, posing a greater challenge in detecting unsolvable states. A similar trend is also observed in other puzzles that require more depth of exploration to solve, including graph coloring and grid puzzles (see Appendix A.4). These observations reveal that general LLMs tend to make overly optimistic decisions (i.e., assuming a solvable state) when faced with tasks that exceed their actual capabilities.

### v2

@@ Abstract
Many challenging reasoning tasks require not just rapid, intuitive responses, but a more deliberate, multi-step approach. Recent progress in large language models (LLMs) highlights an important shift from the “System 1” way of quick reactions to the “System 2” style of reflection-and-correction problem solving. However, current benchmarks heavily rely on the final-answer accuracy, leaving much of a model’s intermediate reasoning steps unexamined. This fails to assess the model’s ability to reflect and rectify mistakes within the reasoning process. To bridge this gap, we introduce FineReason, a logic-puzzle benchmark for systematic evaluation of LLMs’ reasoning capabilities. Each puzzle can be decomposed into atomic steps, making it ideal for rigorous validation of intermediate correctness. Building on this, we introduce two tasks: state checking and state transition, for a comprehensive evaluation of how models assess the current situation and plan the next move. To support broader research, we also provide a puzzle training set aimed at enhancing general reasoning. We show that models trained on our state checking and transition data demonstrate gains in mathematical reasoning by up to $5.1\%$ [^1] https://github.com/DAMO-NLP-SG/FineReason
@@
Guizhen Chen1,2,thanks: Guizhen and Chaoqun are under the Joint PhD Program between Alibaba and NTU.  Weiwen Xu2,thanks: Corresponding authors.  Hao Zhang2,22footnotemark: 2  Hou Pong Chan2 Chaoqun Liu1,2,11footnotemark: 1Lidong Bing2Deli Zhao2,3Anh Tuan Luu1Yu Rong2,3 1 Nanyang Technological University, Singapore 2 DAMO Academy, Alibaba Group, Singapore 3 Hupan Lab, Hangzhou, China
@@ 1 Introduction
Figure 1: A simplified Sudoku solution tree illustrating stepwise state checking (solvability prediction) and transition (next move determination, including forward moves and backtracking).
Figure 2: An illustration of four puzzle categories in FineReason. These puzzles are solved through discrete steps, with explicit rules allowing easy validation of intermediate states.
@@
In cognitive science, human reasoning is typically characterized by two distinct systems: (i) System 1, which is fast, automatic, and effortless, and (ii) System 2, which is slow, analytical, and effortful (Kahneman, 2011). System 2 reasoning enables humans to proactively anticipate future outcomes, reassess intermediate states, and iteratively refine solutions Yao et al. (2023), thereby allowing them to tackle more complex reasoning tasks. Recent studies suggest that large language models (LLMs) can attain advantages akin to those of System 2 reasoning OpenAI (2024); Snell et al. (2025); DeepSeek-AI et al. (2025); Team (2025). Instead of generating answers directly, these models can iteratively reflect and correct their reasoning Shinn et al. (2023), achieving strong performance on reasoning-intensive tasks like mathematics and coding Qin et al. (2024); Li et al. (2024a); Muennighoff et al. (2025).
@@
Despite these improvements, the underlying mechanisms remain underexplored. Existing reasoning benchmarks primarily focus on the final-answer accuracy Hendrycks et al. (2021a); Cobbe et al. (2021); Chen et al. (2021); Hendrycks et al. (2021b), which offers limited insight into LLMs’ internal reasoning processes, such as reflection and correction. For instance, models might reach a correct conclusion through flawed reasoning Zelikman et al. (2022); Creswell et al. (2023); Lightman et al. (2024); Chia et al. (2024a). This diminishes the trustworthiness of model outputs, posing potential risks in practical usage. Moreover, models might achieve high accuracy by exploiting superficial patterns in the training data Roelofs et al. (2019); Xu et al. (2023); Enström et al. (2024), making it difficult to ascertain whether the observed performance gain truly stems from deliberate reasoning. Therefore, there is a pressing need for more comprehensive reasoning benchmarks that assess the integrity of intermediate processes.
@@
In this work, we present FineReason, a logic-puzzle benchmark for evaluating LLMs’ deliberate reasoning capabilities, such as reflection, backtracking, and exploring alternative solutions. FineReason includes four types of puzzles: Sudoku, Graph Coloring, Game of 24, and Grid Puzzles, as shown in Figure 2. Solving a logic puzzle involves a series of discrete steps, and the explicit rules make it straightforward to validate the intermediate states. To structure our evaluation, we propose two key actions in each atomic step, state checking and state transition, as illustrated in Figure 1. State checking predicts whether the current state can lead to a solvable solution Agarwal et al. (2019); Wang et al. (2024). It captures both retrospective evaluation of prior steps Lightman et al. (2024) and prospective analysis of future steps. Meanwhile, state transition focuses on determining the next valid step, either moving forward or backtracking to the previous state. Together, these two tasks cover the entire puzzle-solving process, revealing the internal reasoning processes of reflection, correction, and exploration of alternative paths in LLMs.
Our evaluation reveals a significant 19.7% performance gap between OpenAI-o1 OpenAI (2024) and Gemini-2.0-Flash-Thinking Google (2024), a difference not captured by other maths and coding benchmarks where high scores saturate. General-purpose models like GPT-4o OpenAI et al. (2024) struggle with deliberate reasoning in FineReason, making near-random guesses in state checking and poor performance in state transition.
To enhance reasoning, we develop a specialized training set on puzzle state checking and transition. Integrating our dataset with common reasoning data consistently improves performance on complex reasoning tasks. For example, when applied to DeepSeek-R1-Distill-Qwen-7B, our puzzle data boosts the accuracy from 82.3% to 87.4% on GSM8K, compared to models trained exclusively on math data. Our results suggest that skills such as backtracking and constraint validation generalize from puzzles to general reasoning, similar to how structured practice (e.g., chess) enhances human strategic thinking and problem solving.
@@
Our main contributions are three-fold: 1) We introduce FineReason, a puzzle-based benchmark accompanied by systematic evaluations on state checking and transition, to provide a more precise evaluation of models’ reasoning abilities, particularly in reflection and correction. 2) Experimental results reveal substantial limitations in deliberate reasoning among general-purpose models, and even in the leading reasoning models. 3) We show that training on structured puzzles transfers deliberate reasoning skills to general problem-solving.
@@ 2 FineReason
We present FineReason, a logic-puzzle benchmark that comprehensively assesses LLMs’ reasoning through stepwise evaluation of state checking and transition. Inspired by the adage “think twice before acting,” these actions capture how models assess the current situation (i.e., state checking) and plan the next move (i.e., state transition), skills crucial for deliberate reasoning.
Formally, the reasoning process of a logic puzzle can be represented as $p=\{p_{1},p_{2},...,p_{n}\}$, where $n$ denotes the number of atomic steps. Each step $p_{i}$ consists of a puzzle state $s_{i}$ and two actions: state checking $a^{c}_{i}$ and state transition $a^{t}_{i}$, i.e., $p_{i}=(s_{i},a^{c}_{i},a^{t}_{i})$. Applying these actions to $s_{i}$ produces the next state $s_{i+1}$. The puzzle-solving process begins at an initial state $s_{1}$ and proceeds through a sequence of atomic steps until reaching the solution state $s_{n}$ that satisfies all constraints.
In the following section, we first introduce a tree-based approach for step decomposition in our puzzles (§2.1). Next, we describe the two key actions – state checking and state transition – that facilitate reasoning evaluation of models (§2.2).
@@ 2.1 Tree-based Puzzle Decomposition
Puzzle solving can be represented as a tree, where nodes are intermediate states, and edges represent the execution of state checking and state transition, as illustrated in Figure 1. Edges are bidirectional, enabling both the exploration of child states and backtracking to the parent when encountering dead ends. This process begins at the root node $s_{1}$ and terminates at a solution leaf $s_{n}$, potentially requiring multiple backtracks to explore different paths.
@@
In Sudoku, the aim is to fill the empty cells such that each row, each column, and each of the $3\times 3$ subgrids contains all digits from 1 to 9 exactly once. A Sudoku state can be either a partially or fully filled $9\times 9$ grid. The minimal move is defined as either adding a number for exploration or removing a number for backtracking. We use Sudoku questions from Kaggle [^2] https://www.kaggle.com/datasets/bryanpark/sudoku to create the Sudoku trees.
@@
The aim of graph coloring is to assign colors to each vertex in a graph such that no two adjacent vertices share the same color. Each puzzle specifies a maximum number of colors allowed in a graph. A graph coloring state is either a partially colored graph or a completely colored graph. A minimal move involves either assigning a color to a vertex or removing a color from a vertex. To create the questions, we generate random graphs and find their respective chromatic numbers using the backtracking algorithm van Beek (2006).
@@
In Game of 24, given four numbers, the task is to apply basic arithmetic operations (addition, subtraction, multiplication, and division) to reach exactly the value of 24. Each number must be used exactly once. Each state is a partial or complete arithmetic expression. The minimal move is to apply or unapply an operation to two of the remaining numbers. We use the problem set from Yao et al. (2023) to generate the trees.
@@ 2.2 Evaluation Tasks
We define two key actions, state checking and state transition, which enable movements between states.
@@
State checking requires the model to assess if a given state $s_{i}$ can lead to a solvable solution $s_{n}$. Based on our constructed puzzle trees, we label a state as solvable if at least one valid solution exists in the subtree of $s_{i}$. To ensure a diverse difficulty range, we uniformly sample both solvable and unsolvable states across different tree depths. For each sampled state, we prompt models to evaluate its solvability with puzzle rules, prior visited states, and the current state (see Appendix A.2). In general, state checking evaluates two key aspects: 1) It tests if models can reflect on prior steps to avoid rule violations (e.g., preventing duplicate “1”s in a Sudoku row). 2) It requires models to anticipate potential dead ends by looking ahead. The second aspect, however, requires a higher level of reasoning to proactively detect unsolvable states.
@@
State transition involves making the minimal move defined in §2.1, which requires models to determine the next valid state. Based on the state-checking results, models should proceed if the state is solvable and backtrack otherwise. Specifically, at a solvable state, a correct transition would be an unvisited child of the given state. At an unsolvable state, the correct move is to backtrack to its parent state. To isolate the impact of state transition from incorrect state checking, our evaluation provides ground-truth state-checking labels. We sample states from the puzzle tree and construct prompts with puzzle rules, prior visited states, the current state, and some unsolvable child states (see Appendix A.2). The inclusion of unsolvable child states is to assess whether models can effectively bypass these bad states.
@@
Table 2: A preliminary study on end-to-end puzzle-solving performance of LLMs.
@@ 3 Experimental Setup
Implementation.
We use zero-shot chain of thought prompting Kojima et al. (2022) for evaluation. To ensure a genuine evaluation of LLM’s inherent reasoning capabilities, we explicitly include “Do not solve using programming” in the instruction, restricting models from relying on memorized code snippets from their training data. In our actual attempts, without explicitly stating this constraint, models tend to generate Sudoku solvers or graph-coloring algorithms instead of demonstrating deliberate reasoning. Prompt templates are shown in Appendix A.2.
@@
We select the best-performing open and closed-source models, including 1) reasoning-oriented models: o1 OpenAI (2024), Gemini-2.0-Flash-Thinking (Gemini-FT, Google (2024)), and 2) general-purpose models that perform straightforward execution: GPT-4o OpenAI et al. (2024), GPT-3.5 OpenAI (2022), Gemini-2.0-Flash (Gemini-F, Google (2024)), and Qwen2.5-72B-Instruct (Qwen2.5-Inst, Qwen et al. (2025)).
@@ 4 Evaluation Results
In this section, we first present a preliminary evaluation of LLMs on end-to-end puzzle solving (§4.1), which reveals inconsistencies in model performance. To gain deeper insights beyond their end-to-end performance, we present our main results focusing on state checking and transition tasks (§4.2), followed by analyses on models’ behaviors, error patterns, and performance across different difficulty levels (§4.3 to §4.6).
@@ 4.1 Preliminary: End-to-End Puzzle Solving
Table 2 presents an initial evaluation of LLMs on end-to-end puzzle-solving tasks. Despite their strong performance on coding and math tasks Qwen et al. (2025), these models are notably weak in end-to-end puzzle solving. Additionally, there are some counter-intuitive observations: Gemini-F outperforms Gemini-FT on Sudoku and Game of 24, yet struggles on the other two puzzles. These inconsistencies suggest that end-to-end puzzle solving alone may not be a reliable metric for assessing LLMs’ reasoning, emphasizing the need for more granular evaluation methods.
@@ 4.2 Main results
To understand models’ reasoning capabilities in greater depth, we evaluate on state checking and transition. Results in Table 3 reveal noticeable performance gaps between reasoning-oriented and general-purpose models. On the state-checking task, reasoning-oriented models (o1 and Gemini-FT) consistently lead the performance in every puzzle category. In contrast, general-purpose models barely match the random baseline in puzzles like Sudoku and Grid Puzzles. A similar trend is observed on the state-transition task. These findings further support the view that inference-time scaling can substantially boost reasoning capabilities Snell et al. (2025); Muennighoff et al. (2025).
Between the reasoning models, Gemini-FT generally performs on par with o1 in state checking but consistently lags behind in state transition. This reveals weaknesses in Gemini-FT’s reasoning process, particularly in error revision. These findings align well with our practical experience using these models, which provides empirical evidence that FineReason offers a more accurate reflection of LLMs’ reasoning capabilities.
4.3 Model Behaviors in State Checking
State checking requires looking ahead to identify unsolvable states. To analyze models’ behaviors, we report the state-checking precision, recall, and F1 scores in Table 4, designating unsolvable states as positive cases. Recall measures the model’s ability to detect dead ends, while precision indicates the reliability of its unsolvable state predictions.
We find that reasoning models generally detect unsolvable states well, as indicated by the high F1 scores. As for general models (GPT-4o, Gemini-F, Qwen2.5-Inst), the recall is generally low in Sudoku but not in Game of 24, possibly due to Sudoku’s deeper puzzle tree making unsolvable state detection harder. A similar trend is also observed in other puzzles that have deeper solutions, including graph coloring and grid puzzles (see Appendix A.3). These observations reveal that general LLMs tend to make overly optimistic decisions (i.e., assuming a solvable state) when faced with tasks that exceed their actual capabilities.